# Mitigating distribution shift in machine learning-augmented hybrid simulation

### v1

@@ Abstract
We study the problem of distribution shift generally arising in machine-learning augmented hybrid simulation, where parts of simulation algorithms are replaced by data-driven surrogates. We first establish a mathematical framework to understand the structure of machine-learning augmented hybrid simulation problems, and the cause and effect of the associated distribution shift. We show correlations between distribution shift and simulation error both numerically and theoretically. Then, we propose a simple methodology based on tangent-space regularized estimator to control the distribution shift, thereby improving the long-term accuracy of the simulation results. In the linear dynamics case, we provide a thorough theoretical analysis to quantify the effectiveness of the proposed method. Moreover, we conduct several numerical experiments, including simulating a partially known reaction-diffusion equation and solving Navier-Stokes equations using the projection method with a data-driven pressure solver. In all cases, we observe marked improvements in simulation accuracy under the proposed method, especially for systems with high degrees of distribution shift, such as those with relatively strong non-linear reaction mechanisms, or flows at large Reynolds numbers.
@@ 1 Introduction
Many scientific computational applications, such as computational fluid dynamics (CFD) and molecular dynamics (MD) can be viewed as dynamical system modeling and simulation problems, which is tackled by rigorous numerical tools with theoretical guarantee [45, 26]. However, in many cases a part of the simulation workflow, such as the Reynolds stresses in Reynolds-averaged Navier-Stokes equation (RANS) [1] and the exchange-correlation energies in density function theory used to compute force-fields that drives MD simulations [27], depends on models that are either expensive to compute, or even unknown in practice. One thus often resort to a hybrid simulation method, where the known, resolved components of the dynamics are computed exactly, while the unresolved components are replaced by approximate, but computationally tractable models. For example, solving Navier-Stokes equations using projection method involves two steps. In the first step, all the terms except for the gradient of the pressure is used to evolve the velocity. This step is computationally cheap and thus understood as the resolved part. Next, the pressure is solved from a Poisson equation and then used to correct the velocity. Most of the computational cost is contained in solving this Poisson equation and we thereby viewed this as the unresolved part. We call such scientific computing problems with both resolved and unresolved parts “hybrid simulation problems”. Similar problems are surveyed in [51] under the name of “Hybrid physics-DL models”.
@@
As machine learning becomes increasingly powerful in areas like computer vision and natural language processing, practitioners begin to use data-driven modules to model the unresolved part to carry out the simulation. For example, in [49] the author replaced the numerical Poisson solver of the unresolved part by a convolutional neural network trained using a novel unsupervised learning framework. Then, this data-driven model is coupled with resolved model and provide fast and realistic simulation results in 2D and 3D. Similar ideas are used else where, e.g. RANS [28]. We hereafter refer to this simulation procedure by machine-learning augmented hybrid simulation (MLHS). This structure represents a great part of the scientific machine learing research [54, 46, 25] and will be the focus of this paper. In MLHS, a common problem appears: while the data-drive model performs well on the training data, the performance quickly deteriorates when iteratively applying it in simulation, driving the dynamics to some regime which is not observed from the training source. Empirical evidences have been observed in various applications, such as CFD [46, 35, 9, 54], molecular dynamics [56, 55, 25], and iterative numerical solver [2] but these problems are not well-studied algorithmically and theoretically in the literature.
@@
This issue has strong connections with the so-called distribution shift (DS) in computer science applications, especially in reinforcement learning and computer vision. Researchers use DS to refer to problems where training and testing distribution differ significantly. An example is an image classifier trained with images taken in the daytime that is tested under night conditions [23]. Therefore, accuracy on the training regime itself cannot guarantee the performance during inference. To resolve this, researchers have systematically developed several methods, e.g. domain adaptation [11], ensemble learning [8], and meta-learning [50]. However, there are key differences between the distribution shift phenomena in MLHS and that in traditional applications between the DS phenomena in MLHS and computer vision. While distribution shift in computer vision lacks a theoretical model to describe [23], that in MLHS comes from dynamical systems for which we have abundant knowledge on the resolved parts of the models. For example, let us consider using projection method [10] to solve the incompressible Navier-Stokes equation. Suppose we replace the Poisson pressure solver by a data-driven model based on neural network. Then, the dynamics of this data-driven hybrid numerical solver will largely depend on the properties of the resolved numerical part of the projection method, which is well-studied [10]. By analyzing the stability properties of this dynamics, one can quantify which family of distribution shifts may arise and resolve them according to the information of the dynamics. Therefore, one can pose the following question: Can we use the information of resolved parts to design robust learning algorithm for unresolved parts to improve hybrid simulation?
@@
In this paper, we first develop a mathematical framework to understand the origin of distribution shift in MLHS and how it may lead to simulation performance deterioration. We emphasize the difference between this instability issue in data-driven scientific computing and distribution shift in the machine learning literature, such as reinforcement learning and computer vision. Then, we propose an algorithm to improve the simulation accuracy by mitigating distribution shift. We assume by manifold hypothesis [34] that the correct trajectories lie on a low-dimensional manifold of the high dimensional space, e.g. the fluid configuration of Navier-Stokes equation lies on the solution manifold of the high-dimensional grid space. The key idea is to combine the physical information of the resolved part of the model, i.e. Navier-Stokes equation and this manifold structure learned from the data to form a regularization term for the data-driven model. Intuitively speaking, this regularizer stablizes the dynamics by preventing it from moving further away the data manifold. Such movements may result in configurations that are either non-physical, or corresponds to different initial/boundary conditions. Therefore, preventing such movements reduces the severity of distribution shift, and the data-driven model will stay relatively accurate during the simulation, which promotes high simulation fidelity in a long time interval. In implementation, we first use an autoencoder (AE) to parameterize the underlying data manifold. After this preprocessing, the AE is combined with the resolved dynamics and plugged into the loss function of the model as a regularization, which prevents the simulation from moving to unseen regimes. One then back-propagates this modified loss function to optimize the data-driven model. The algorithm is tested on several representative numerical examples to demonstrate its effectiveness. We show that the proposed approach can improve simulation accuracy under different extents of distribution shift. Indeed, the improvements become more significant in scenarios where the fidelity of the simulation is highly sensitive to errors introduced in the data-driven surrogates, such as fluid simulations at high Reynolds numbers and reaction-diffusion equations with relatively strong non-linear reaction. Specifically, in high Reynolds number fluid simulations, naive data-driven surrogate models may quickly cause error blow-ups, but our method can maintain simulation accuracy over large time intervals.
@@
The paper is structured as follows. In Section 2, we establish a precise framework to identify the origins of distribution shift in MLHS. In Section 3, we introduce our regularized learning algorithm motivated from the analysis of the previously identified form of distribution shift and theoretically understand its performance in the linear setting. We also discuss connections with the literature on control and system identification. In Section 4, we validate our algorithm on several practical numerical cases, including simulating a reaction-diffusion equation and the incompressible Navier-Stokes equation.
@@ 2 Mathematical formulation of MLHS and the problem of distribution shift
In this section, we provide the mathematical formulation of MLHS and then identify the problem of distribution shift. In the first subsection we give a general treatment, and we provide concrete examples in Section 2.2.
@@ 2.1 The resolved and unresolved components in hybrid simulations
$\partial_{t}\mathbf{u}$ $=\mathcal{L}(\mathbf{u},\mathbf{y},t),\quad\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},\mathcal{L}:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{m},$ (1)
$\mathbf{y}$ $=\phi(\mathbf{u},t),\quad\phi:\mathbb{R}^{m}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{n},$
where $\mathbf{u}$ is the resolved state variable, e.g. fluid velocity field or chemical concentration field. The vector $\mathbf{y}$ is the unresolved variable, which drives the dynamics for $\mathbf{u}$ but is either expensive to compute or cannot be directly observed. Throughout the paper, we will omit writing the explicit time-dependence by using $\mathbf{u}$ ($\mathbf{y}$) to denote $\mathbf{u}(t)$ ($\mathbf{y}(t)$) whenever there is no ambiguity. We also restrict the discussion to the finite dimension case, but most of the analysis here will be applicable to the more general case of $\mathbf{u},\mathbf{y}$ belonging to an infinite dimensional Hilbert space. We adopt the following assumptions associated with Eq. 1
@@
  • 2. Unresolved component: $\phi$ is either unknown, partially-known or expensive to evaluate.
Given such a system, the goal is to first obtain information on the unresolved model and then integrate it with resolved part to simulate the whole dynamics. This hybrid structure, which we call hybrid simulation problems is general enough to include most of the settings in hybrid simulation, e.g. when simulating the trajectories of molecular dynamics which satisfy some stochastic differential equations or the time-evolution of fluid velocity fields following the Navier-Stokes equation. This paper focuses on MLHS, a particular variation of such hybrid simulation problem where the unresolved component is tackled by a data-driven method. This point will be made more precise in Section 2.3. To simplify the analysis, we adopt the forward Euler time discretization for simulation
$\widehat{\mathbf{u}}_{k+1}$ $=\widehat{\mathbf{u}}_{k}+\Delta t\mathcal{L}(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k}),\qquad\widehat{\mathbf{y}}_{k}=\phi(\widehat{\mathbf{u}}_{k}),$ (2)
while other consistent discretizations may be analyzed in the same spirit. Here we drop the dependence of $\mathcal{L},\phi$ on $t$ to consider autonomous system while our later discussion and algorithm are also applicable to non-autonomous system.
@@ 2.2 Examples of hybrid simulation problems
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nu\Delta\mathbf{u}=\nabla p,\quad\nabla\cdot\mathbf{u}=0,\quad T\in[0,1],$ (3)
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ is velocity and $p$ pressure. Fix a regular grid size for discretization. The projection method can be written as follows
$\mathbf{u}_{k+1}$ $=\mathbf{u}_{k}+\Delta t\mathcal{L}(\mathbf{u}_{k},p_{k})$ (4)
$=\mathbf{u}_{k}+\Delta t(\nu\Delta\mathbf{u}_{k}-(\mathbf{u}_{k}\cdot\nabla)\mathbf{u}_{k}-\nabla p_{k}),$
@@
We write $\Delta^{-1}$ as the inverse operator for Poisson equation formally. The resolved part is performed by first stepping the convection and diffusion term then using pressure to correct the step, and the unresolved part is a Poisson equation which related the unsolved state $p_{k}$ with the velocity $\mathbf{u}_{k}$. Here, $p_{k}$ represents the unresolved variable $\mathbf{y}_{k}$ in our formulation Eq. 1.
In this method, the most expensive step is the pressure computation, which requires repeated solutions of similar large-scale linear equations. State of the art solvers such as multigrid [4] and conjugate gradient [43] become prohibitively expensive when the problem size is large. Recently, a promising direction is to replace the Poisson solver by data-driven surrogate models. Ref. [49] replaces the pressure calculation step by a convolutional neural network trained with unsupervised learning by requiring the updated velocity to have zero divergence. Similarly, machine learning models such as tensor-based neural networks are used in [28] to replace classical turbulence modeling for unclosed Reynolds stresses tensor. This model is then plugged into the Reynolds-averaged Navier-Stokes simulation to predict the flow separation. More recently, [37] combines the idea of reduced-order modeling with graph convolutional neural network to encode the reduced manifold and enable fast evaluations of parametrized PDEs.
The second example is based on simulation on two grids of different sizes. Let us continue with the projection method setting Eq. 4. Denote the time-evolution operator on grid size $n$ as $f_{n}:\mathbf{u}_{k}^{n}\rightarrow\mathbf{u}_{k+1}^{n}$, which becomes very expensive to compute when $n$ is large. Therefore, one may wish to replace the fine-grid solver with a coarse-grid one and add some correction. This again reduces the computational burden for high-fidelity simulations. This structure shows up in various scientific computing situations, such as multigrid solvers [4], Reynolds-averaged Navier-Stokes (RANS) [38, 1], and large eddy simulation (LES) [59]. Suppose we have a fine grid with size $2n\times 2n$ and a coarse grid with size $n\times n$, we use superscript $n,2n$ to denote the field defined on these two grids, i.e. $\mathbf{u}_{k}^{n},\mathbf{u}_{k}^{2n}$. Moreover, we fix an interpolation and a restriction operator between fields on these two grids:
$I_{n}^{2n}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{2n\times 2n},\quad R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n}.$ (5)
@@
In detail, the resolved component contains a solver of the evolution equation over the coarse grid which marches forward by one time step, with field given by a restriction of the field on the fine grid. Then, the next step field configuration is interpolated to the fine grid, and the unresolved component serves to correct the deviation of the field variables between simulating on grids of different sizes. The unresolved variables $\mathbf{y}_{k}^{2n}$ can be calculated if we have accurate simulation results $\mathbf{u}_{k}^{2n}$ on the fine grid by
$\mathbf{y}_{k}^{2n}=\mathbf{u}_{k+1}^{2n}-I_{n}^{2n}\circ f_{n}(R_{2n}^{n}(\mathbf{u}_{k}^{2n}))).$ (7)
In the literature, [21] treats the smoothing algorithm in multigrid solver as unresolved components and other procedures such as restriction and prolongation as resolved components. This fits into our framework of MLHS. They introduce a supervised loss function based on multigrid convergence theory to learn the optimal smoother and improve the convergence rate over anisotropic rotated Laplacian problems and variable coefficient diffusion problems. In [32], the authors applies neural network as non-linear regression to fit between some key fluid features and sub-grid stresses. Going beyond this, [35] adds neural emulation to offline learning to prevent the trajectories from deviating away the ground truth. This data-driven model is plugged into hybrid simulation to test its ability of preserving coherent structure and scaling laws.

### v2

@@ Abstract
We study the problem of distribution shift generally arising in machine-learning augmented hybrid simulation, where parts of simulation algorithms are replaced by data-driven surrogates. A mathematical framework is established to understand the structure of machine-learning augmented hybrid simulation problems and the cause and effect of the associated distribution shift. We show correlations between distribution shift and simulation error both numerically and theoretically. Then, we propose a simple methodology based on tangent-space regularized estimator to control the distribution shift, thereby improving the long-term accuracy of the simulation results. In the linear dynamics case, we provide a thorough theoretical analysis to quantify the effectiveness of the proposed method. Moreover, we conduct several numerical experiments, including simulating a partially known reaction-diffusion equation and solving Navier-Stokes equations using the projection method with a data-driven pressure solver. In all cases, we observe marked improvements in simulation accuracy under the proposed method, especially for systems with high degrees of distribution shift, such as those with relatively strong non-linear reaction mechanisms, or flows at large Reynolds numbers.
@@ 1 Introduction
Many scientific computational applications, such as computational fluid dynamics (CFD) and molecular dynamics (MD) can be viewed as dynamical system modeling and simulation problems, which are tackled by rigorous numerical tools with theoretical guarantee [53]. However, in many cases a part of the simulation workflow, such as the Reynolds stresses in Reynolds-averaged Navier-Stokes equation (RANS) [1] and the exchange-correlation energies in density function theory used to compute force fields that drive MD simulations [30], depends on models that are either expensive to compute or even unknown in practice. One thus often resorts to a hybrid simulation method, where the known, resolved components of the dynamics are computed exactly, while the unresolved components are replaced by approximate, but computationally tractable models. For example, solving Navier-Stokes equations using the projection method involves two steps. In the first step, all the terms except for the gradient of the pressure are used to evolve the velocity. This step is computationally cheap and thus understood as the resolved part. Next, the pressure is solved from a Poisson equation and then used to correct the velocity. Most of the computational cost is contained in solving this Poisson equation, and we thereby viewed this as the unresolved part. We call such scientific computing problems with both resolved and unresolved parts “hybrid simulation problems”. Similar problems are surveyed in [59] under the name of “Hybrid physics-DL models”.
@@
As machine learning becomes increasingly powerful in areas like computer vision and natural language processing, practitioners begin to use data-driven modules to model the unresolved part to carry out the simulation. For example, in [57] the author replaced the numerical Poisson solver of the unresolved part with a convolutional neural network trained using a novel unsupervised learning framework. Then, this data-driven model is coupled with resolved models and provides fast and realistic simulation results in 2D and 3D. Similar ideas are used elsewhere, e.g. RANS [31] and LES [50]. We hereafter refer to this simulation procedure as machine-learning augmented hybrid simulation (MLHS). This structure represents a great part of the scientific machine learning research [62, 54] and will be the focus of this paper. In MLHS, a common problem appears: while the data-drive model performs well on the training data, the performance quickly deteriorates when iteratively applying it in simulation, driving the dynamics to some regimes that are not observed from the training source. Empirical evidences have been observed in various applications, such as CFD [50, 54, 39, 62], molecular dynamics [64, 63, 29], and iterative numerical solver [2], but these problems are not well-studied algorithmically and theoretically in the literature.
@@
This issue has strong connections with the so-called distribution shift (DS) in computer science applications, especially in reinforcement learning and computer vision. Researchers use DS to refer to problems where training and testing distribution differ significantly. An example is an image classifier trained with images taken in the daytime that is tested under night conditions [25]. Therefore, the accuracy during the training itself cannot guarantee the performance during inference. To resolve this, researchers have systematically developed several methods, e.g. domain adaptation [10], ensemble learning [7], and meta-learning [58]. However, there are key differences between the distribution shift phenomena in MLHS and that in traditional applications between the DS phenomena in MLHS and computer vision. While distribution shift in computer vision lacks a theoretical model to describe [25], that in MLHS comes from dynamical systems for which we have abundant knowledge on the resolved parts of the models. For example, let us consider using the projection method [8] to solve the incompressible Navier-Stokes equation. Suppose we replace the Poisson pressure solver with a data-driven model based on neural networks. Then, the dynamics of this data-driven hybrid numerical solver will largely depend on the properties of the resolved numerical part of the projection method, which is well-studied [8]. By analyzing the stability properties of the dynamics, one can quantify which family of distribution shifts may arise and resolve them according to the information of the dynamics. Therefore, one can pose the following question: Can we use the information of resolved parts to design robust learning algorithms for unresolved parts to improve hybrid simulation?
@@
In this paper, we first develop a mathematical framework to understand the origin of distribution shift in MLHS and how it may lead to simulation performance deterioration. We emphasize the difference between this instability issue in data-driven scientific computing and distribution shift in the machine learning literature, such as reinforcement learning and computer vision. Then, we propose an algorithm to improve the simulation accuracy by mitigating distribution shifts. We assume by manifold hypothesis [37] that the correct trajectories lie on a low-dimensional manifold of the high-dimensional space, e.g. the fluid configuration of the Navier-Stokes equation lies on the solution manifold of the high-dimensional grid space. The key idea is to combine the physical information of the resolved part of the model, i.e. Navier-Stokes equation and this manifold structure learned from the data to form a regularization term for the data-driven model. Intuitively speaking, this regularizer stabilizes the dynamics by preventing it from moving further away from the data manifold. Such movements may result in configurations that are either non-physical, or corresponds to different initial/boundary conditions. Therefore, preventing such movements reduces the severity of distribution shift, and the data-driven model will stay relatively accurate during the simulation, which promotes high simulation fidelity in a long time interval. In implementation, we first use an autoencoder (AE) to parameterize the underlying data manifold. After this preprocessing, the AE is combined with the resolved dynamics and plugged into the loss function of the model as a regularization, which prevents the simulation from moving to unseen regimes. One then back-propagates this modified loss function to optimize the data-driven model. The algorithm is tested on several representative numerical examples to demonstrate its effectiveness. We show that the proposed approach can improve simulation accuracy under different extents of distribution shift. Indeed, the improvements become more significant in scenarios where the fidelity of the simulation is highly sensitive to errors introduced in the data-driven surrogates, such as transport-dominant fluid simulations and reaction-diffusion equations with relatively strong non-linear reactions. Specifically, in fluid simulations, naive data-driven surrogate models may quickly cause error blow-ups, but our method can maintain simulation accuracy over large time intervals.
@@
The paper is structured as follows. In Section 2, we establish a precise framework to identify the origins of distribution shift in MLHS. In Section 3, we introduce our regularized learning algorithm motivated by the analysis of the previously identified form of distribution shift and theoretically understand its performance in the linear setting. We also discuss connections with the literature on control and system identification. In Section 4, we validate our algorithm on several practical numerical cases, including simulating a reaction-diffusion equation and the incompressible Navier-Stokes equation.
@@ 2 Mathematical formulation of MLHS and the problem of distribution shift
In this section, we provide the mathematical formulation of MLHS and then identify the problem of distribution shift. In the first subsection, we give a general treatment, and we provide concrete examples in Section 2.2.
@@ 2.1 The resolved and unresolved components in hybrid simulations
(1) $\partial_{t}\mathbf{u}=\mathcal{L}(\mathbf{u},\mathbf{y},t),\quad\mathbf{y}=\phi(\mathbf{u},t),$
where $\mathbf{u}$ is the resolved state variable, e.g. fluid velocity field or chemical concentration field. The vector $\mathbf{y}$ is the unresolved variable that drives the dynamics for $\mathbf{u}$ but is either expensive to compute or cannot be directly observed. Throughout the paper, we will omit writing the explicit time-dependence by using $\mathbf{u}$ ($\mathbf{y}$) to denote $\mathbf{u}(t)$ ($\mathbf{y}(t)$) whenever there is no ambiguity. In numerical simulation, one always discretizes the ODE or PDE to obtain a finite-dimensional system, i.e. $\mathbf{u}\in\mathbb{R}^{m},\mathbf{y}\in\mathbb{R}^{n},\mathcal{L}:\mathbb{R}^{m}\times\mathbb{R}^{n}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{m},\phi:\mathbb{R}^{m}\times\mathbb{R}_{+}\rightarrow\mathbb{R}^{n}$. Therefore, we restrict the discussion to the finite dimension case where both $\mathcal{L},\phi$ becomes mapping between finite-dimensional spaces, but most of the analysis here applies to the more general case of $\mathbf{u},\mathbf{y}$ belonging to an infinite dimensional Hilbert space. We adopt the following assumptions associated with Eq. 1:
@@
  • 2. Unresolved component: $\phi$ is either unknown or expensive to evaluate.
Given such a system, the goal is to first obtain information on the unresolved model and then integrate it with the resolved part to simulate the whole dynamics. This hybrid structure, which we call hybrid simulation problems, is general enough to include many applications, e.g. when simulating the trajectories of molecular dynamics which satisfy some stochastic differential equations or the time-evolution of fluid velocity fields following the Navier-Stokes equation. We focus on MLHS, a particular variation of such hybrid simulation problems where the unresolved component is tackled by a data-driven method. This point will be made more precise in Section 2.3. To simplify the analysis, we adopt the forward Euler time discretization for simulation
(2) $\widehat{\mathbf{u}}_{k+1}$ $=\widehat{\mathbf{u}}_{k}+L(\widehat{\mathbf{u}}_{k},\widehat{\mathbf{y}}_{k})\Delta t,\qquad\widehat{\mathbf{y}}_{k}=\phi(\widehat{\mathbf{u}}_{k}),$
while other consistent discretizations may be analyzed in the same spirit. Here we drop the dependence of $L,\phi$ on $t$ to consider autonomous systems while our later discussion and algorithm are also applicable to non-autonomous systems.
@@ 2.2 Examples of hybrid simulation problems
(3) $\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nu\Delta\mathbf{u}=\nabla p,\quad\nabla\cdot\mathbf{u}=0,\quad T\in[0,1],$
where $\mathbf{u}=(u(x,y,t),v(x,y,t))^{T}\in\mathbb{R}^{2}$ is the velocity and $p$ is the pressure. Fix a regular grid size for discretization. The projection method can be written as follows
(4) $\mathbf{u}_{k+1}$ $=\mathbf{u}_{k}+L(\mathbf{u}_{k},p_{k})\Delta t=\mathbf{u}_{k}+(\nu\Delta\mathbf{u}_{k}-(\mathbf{u}_{k}\cdot\nabla)\mathbf{u}_{k}-\nabla p_{k})\Delta t,$
@@
We write $\Delta^{-1}$ as the formal inverse operator for the Laplace operator. The resolved part is performed by first stepping the convection and diffusion term then using pressure to correct the step, and the unresolved part is a Poisson equation that relates the unsolved state $p_{k}$ with the velocity $\mathbf{u}_{k}$. Here, $p_{k}$ represents the unresolved variable $\mathbf{y}_{k}$ in our formulation Eq. 1.
In this method, the most expensive step is the pressure computation, which requires repeated solutions of similar large-scale linear equations. State-of-the-art solvers such as multigrid [5] and conjugate gradient [51] become prohibitively expensive when the problem size is large. Recently, a promising direction is replacing the Poisson solver by data-driven surrogate models. Ref. [57] replaces the pressure calculation step with a convolutional neural network trained with unsupervised learning by requiring the updated velocity to have zero divergence. In the same spirit, machine learning models such as tensor-based neural networks are used in [31] to replace classical turbulence modeling for unclosed Reynolds stress tensors. This model is then plugged into the Reynolds-averaged Navier-Stokes simulation to predict the flow separation. More recently, [41] combines the idea of reduced-order modeling with graph convolutional neural network to encode the reduced manifold and enable fast evaluations of parameterized PDEs.
The second example concerns simulations on two grids of different sizes. Let us continue with the projection method setting Eq. 4. Denote the time-evolution operator on grid size $n$ as $f_{n}:\mathbf{u}_{k}^{n}\rightarrow\mathbf{u}_{k+1}^{n}$, which becomes very expensive to compute when $n$ is large. Therefore, one may wish to replace the fine-grid solver with a coarse-grid one and add some corrections. This again reduces the computational burden for high-fidelity simulations. This structure shows up in various scientific computing situations, such as multigrid solvers [5], Reynolds-averaged Navier-Stokes (RANS) [42, 1], and large eddy simulation (LES) [67]. Suppose we have a fine grid with size $2n\times 2n$ and a coarse grid with size $n\times n$, we use superscript $n,2n$ to denote the field defined on these two grids, i.e. $\mathbf{u}_{k}^{n},\mathbf{u}_{k}^{2n}$. Moreover, we fix an interpolation and a restriction operator between fields on these two grids:
(5) $I_{n}^{2n}:\mathbb{R}^{n\times n}\rightarrow\mathbb{R}^{2n\times 2n},\quad R_{2n}^{n}:\mathbb{R}^{2n\times 2n}\rightarrow\mathbb{R}^{n\times n}.$
@@
In detail, the resolved component contains a solver of the evolution equation over the coarse grid which marches forward by one time step, with the field given by a restriction of the field on the fine grid. Then, the next step field configuration is interpolated to the fine grid and the unresolved component serves to correct the deviation of the field variables between simulating on grids of different sizes. The unresolved variables $\mathbf{y}_{k}^{2n}$ can be calculated if we have accurate simulation results $\mathbf{u}_{k}^{2n}$ on the fine grid by $\mathbf{y}_{k}^{2n}=\mathbf{u}_{k+1}^{2n}-I_{n}^{2n}\circ f_{n}(R_{2n}^{n}(\mathbf{u}_{k}^{2n}))$. In the literature, [22] treats the smoothing algorithm in multigrid solver as unresolved components and other procedures such as restriction and prolongation as resolved components. This fits into our framework of MLHS. They introduce a supervised loss function based on multigrid convergence theory to learn the optimal smoother and improve the convergence rate over anisotropic rotated Laplacian problems and variable coefficient diffusion problems. In  [35], the authors apply neural networks with non-linear regression to fit key fluid features and sub-grid stresses. Going beyond this,  [39] adds neural emulation to offline learning to prevent the trajectories from deviating away from the ground truth. This data-driven model is plugged into hybrid simulations to test its ability to preserve coherent structure and scaling laws.